# Gated Tree Cross-attention for Checkpoint-Compatible Syntax Injection in Decoder-Only LLMs

### v1

@@ 5.2 Tree quality and structure-faithfulness controls
Table 6 tests whether performance depends on a linguistically meaningful structure and the Strong Parser column corresponds to GTCA. For the Weak Parser setting, we used the Berkeley parser variant with the smallest parameterization, while the remaining settings are obtained by degrading the strong parses by corrupting the tree structure (e.g., randomizing or permuting the resulting trees). As shown in the table, models benefit only when the parse trees are high quality and coherent. Replacing a strong parser with weaker ones consistently lowers both syntactic generalization and broad benchmark performance, indicating that low quality trees act more like noise than guidance. The corrupt structure further harms performance, with random trees causing a clear drop and permuted trees producing the most severe drop. These results support a structure-faithfulness interpretation that gains arise from consistent hierarchical signals in the parse tree memory rather than from an unconstrained auxiliary attention pathway.
@@ 6 Conclusion
This paper presents a checkpoint-compatible framework for strengthening syntactic competence in decoder-only LLMs by attaching a checkpoint-compatible gated tree cross-attention pathway (GTCA). The approach exposes hierarchical constituency structure through precomputed parse tree memory aligned to token spans, and modulates the cross-attention update with a learned headwise gate, while keeping the backbone architecture unchanged. To limit interference during continued training, the method combines token update masking in MCQA inputs with a three-stage training pipeline that separates task adaptation, structure specialization, and joint refinement. Empirically, GTCA improves syntactic generalization while preserving broad competence across multi evaluations, and this trend holds across different backbones, including Qwen-2.5-7B and Llama-3-8B. Analysis further indicates that gating, token update masking and staged training are important for balancing syntactic gains and retention, and that performance is sensitive to tree quality and coherence, supporting a structure-faithfulness interpretation rather than a pure extra capacity effect.
@@
Appendix A Model Size and Compute
A.1 Model Variants and Parameter Counts
For computational efficiency, the reported results for HellaSwag and Winogrande are evaluated on the validation split. And Table 7 reports the parameter counts and key configuration differences for each variant.
@@ A.1 Model Variants and Parameter Counts
Table 7: Model variants and sizes.
A.2 Overall Compute Budget and computing infrastructure
@@
Appendix B Prompts Used for benchmark
@@
B.1 CLOTH (0-shot) MMLU (0-shot)
@@
B.2 BLiMP (minimal pairs, pairwise preference)
@@
B.3 CoLA (0-shot, binary classification as MCQA)
@@ B.5 Winogrande (5-shot)
Appendix C Grid Search for the structural coefficient $\alpha_{\text{struct}}$
GTCA injects parse tree information through a gated tree cross-attention branch and integrates the resulting structural update via a token update masked weighted by a structural coefficient $\alpha_{\text{struct}}$. In practice, $\alpha_{\text{struct}}$ controls the strength of structure injection and therefore mediates a trade-off between (i) syntactic specialization and (ii) retention of general-purpose capability. We select the target $\alpha_{\text{struct}}$ via grid search, as stated in the main experiments section.
We grid search over a discrete set of target values, a recommended starting grid is: $\alpha_{\text{struct}}^{*}\in\{0.00,0.05,0.10,0.15,0.20\}$. For each candidate $\alpha_{\text{struct}}^{*}$, we use the same staged training pipeline as Table 8 in Section3.4, with:
@@ Appendix C Grid Search for the structural coefficient $\alpha_{\text{struct}}$
  • Stage 2: linearly warm up $\alpha$ from 0 to $\alpha_{\text{struct}}^{*}$ over the recommended list.
@@
We select $\alpha_{\text{struct}}^{*}$ using held-out validation performance under a retention-aware criterion. Table 8 reports the validation results across candidate $\alpha_{\text{struct}}$ values.
@@
Table 8: Grid search over $\alpha$ on the validation split.

### v2

@@ 5.2 Tree Quality and Structure-Faithfulness Controls
Table 6 tests whether performance depends on a linguistically meaningful structure and the Strong Parser column corresponds to GTCA. For the Weak Parser setting, we used the Berkeley parser variant with the smallest parameterization, while the remaining settings are obtained by degrading the strong parses by corrupting the tree structure (e.g., randomizing or permuting the resulting trees). As shown in the table, models benefit only when the parse trees are high quality and coherent. Replacing a strong parser with weaker ones consistently lowers both syntactic generalization and broad benchmark performance, indicating that low quality trees act more like noise than guidance. The corrupt structure further harms performance, with random trees causing a clear drop and permuted trees producing the most severe drop. These results support a structure-faithfulness interpretation that gains arise from consistent hierarchical signals in the parse tree memory rather than from an unconstrained auxiliary attention pathway. This interpretation is further supported by additional robustness analyses in Appendix F, where GTCA remains more stable under controlled grammatical perturbations and mild surface noise than the continued training baselines.
@@ 6 Conclusion
This paper presents a checkpoint-compatible framework for strengthening syntactic competence in decoder-only LLMs by attaching a checkpoint-compatible gated tree cross-attention pathway (GTCA). The approach exposes hierarchical constituency structure through precomputed parse tree memory aligned to token spans, and modulates the cross-attention update with a learned headwise gate, while keeping the backbone architecture unchanged. To limit interference during continued training, the method combines token update masking in MCQA inputs with a three-stage training pipeline that separates task adaptation, structure specialization, and joint refinement. Empirically, GTCA improves syntactic generalization while preserving broad competence across multiple evaluations, and this trend holds across different backbones, including Qwen-2.5-7B and Llama-3-8B. Analysis further indicates that gating, token update masking and staged training are important for balancing syntactic gains and retention, and that performance is sensitive to tree quality and coherence, supporting a structure-faithfulness interpretation rather than a pure extra capacity effect.
@@
Ethics Statement
We use only publicly available pretrained checkpoints and benchmark datasets and follow their licenses and terms of use. No new human participants or manual annotators were involved.
Acknowledgment
The authors would like to thank the anonymous reviewers for their helpful suggestions and comments. This work was supported by the Brain Science and Brain-like Intelligence Technology - National Science and Technology Major Project 2021ZD0204100 (2021ZD0204105 to N. D.).
Appendix A Additional Ablations
We ran a targeted ablation set on the tree encoder and memory construction. Then varied one factor at a time, and keep the rest identical to GTCA. The metrics are in Appendix Table 1. All results are on Qwen-2.5-7B, and we set temperature=0 and disabled sampling (do_sample=False).
The results show that a shared projection across heights reduces all metrics, this supports the idea that different granularities need different parameterization. Learned attention pooling gives a small gain on CoLA, K equals 128 chunks per height gives a small gain on BLiMP, but it adds parameters or introduces another learned attention module. Considering the joint trade-off between resource consumption and throughput we use GTCA default as in the main text.
Setting CLOTH MMLU BLiMP CoLA
GTCA default 83.98 71.02 83.12 56.59
Single shared projection across heights 81.70 70.76 82.08 56.90
Span pooling learned attention pooling 82.04 70.88 80.41 56.04
32 chunks per height 81.25 70.90 82.45 56.20
128 chunks per height 80.36 70.70 83.20 55.05
Appendix Table 1: Ablation results on tree encoder and memory construction. All results are on Qwen-2.5-7B with temperature=0 and do_sample=False.
Appendix B Model Size and Compute
B.1 Model Variants and Parameter Counts
For computational efficiency, the reported results for HellaSwag and WinoGrande are evaluated on the validation split. And Appendix Table 2 reports the parameter counts and key configuration differences for each variant.
@@ B.1 Model Variants and Parameter Counts
Appendix Table 2: Model variants and sizes.
B.2 Overall Compute Budget and Computing Infrastructure
@@
Appendix C Prompts Used for Benchmark
@@
C.1 CLOTH (0-shot) and MMLU (0-shot)
@@
C.2 BLiMP (minimal pairs, pairwise preference)
@@
C.3 CoLA (0-shot, binary classification as MCQA)
@@ C.5 WinoGrande (5-shot)
Appendix D Grid Search for the Structural Coefficient $\alpha_{\text{struct}}$
GTCA injects parse tree information through a gated tree cross-attention branch and integrates the resulting structural update via a token-update-masked residual weighted by a structural coefficient $\alpha_{\text{struct}}$. In practice, $\alpha_{\text{struct}}$ controls the strength of structure injection and therefore mediates a trade-off between (i) syntactic specialization and (ii) retention of general-purpose capability. We select the target $\alpha_{\text{struct}}$ via grid search, as stated in the main experiments section.
We grid search over a discrete set of target values, a recommended starting grid is: $\alpha_{\text{struct}}^{*}\in\{0.00,0.05,0.10,0.15,0.20\}$. For each candidate $\alpha_{\text{struct}}^{*}$, we use the same staged training pipeline as Appendix Table 3 in Section3.4, with:
@@ Appendix D Grid Search for the Structural Coefficient $\alpha_{\text{struct}}$
  • Stage 2: linearly warm up $\alpha$ from 0 to $\alpha_{\text{struct}}^{*}$ over the warm-up steps.
@@
We select $\alpha_{\text{struct}}^{*}$ using held-out validation performance under a retention-aware criterion. Appendix Table 3 reports the validation results across candidate $\alpha_{\text{struct}}$ values.
@@
Appendix Table 3: Grid search over $\alpha$ on the validation split. All results are on Qwen-2.5-7B, and we set temperature=0 and disabled sampling (do_sample=False).
@@
Appendix E Error Analysis
Across the BLiMP phenomena that directly probe word order and movement, the residual errors concentrate in filler-gap dependencies and island effects. These cases are informative because the lexical content is largely preserved while the syntactic configuration changes. The examples below show that the model can still over-rely on local surface compatibility and can underweight global constraints that require tracking long-distance dependencies or respecting extraction islands. This points to clear directions for improvement, such as training objectives that better enforce global constraint satisfaction and evaluation protocols that stress paired stability under controlled syntactic transformations.
Examples of residual errors on filler-gap dependencies and island effects.
For each phenomenon, we show a grammatical sentence, an ungrammatical sentence with a minimal structural change, and the model’s preference. The model consistently prefers the ungrammatical version, indicating overreliance on local surface compatibility and underweighting of global syntactic constraints.
  1. 1.
    Filler-gap dependency
    Grammatical: Kendra saw some ice cream that Maria isn’t hiding.
    Ungrammatical: Kendra saw what Maria isn’t hiding some ice cream.
    Model preference: prefers ungrammatical.
  2. 2.
    Filler-gap dependency
    Grammatical: Gina had known that person that Dan hadn’t forgotten that had criticized many dancers.
    Ungrammatical: Gina had known who that person that Dan hadn’t forgotten had criticized many dancers.
    Model preference: prefers ungrammatical.
  3. 3.
    Filler-gap dependency
    Grammatical: Timothy can conceal who those guys that come here weren’t scaring.
    Ungrammatical: Timothy can conceal that those guys that come here weren’t scaring.
    Model preference: prefers ungrammatical.
  4. 4.
    Island effects
    Grammatical: Who wouldn’t Mitchell investigate he is kissing?
    Ungrammatical: Who wouldn’t Mitchell investigate who is kissing?
    Model preference: prefers ungrammatical.
  5. 5.
    Island effects
    Grammatical: Who would actresses’ hiding most guests stun.
    Ungrammatical: Who would actresses’ hiding stun most guests.
    Model preference: prefers ungrammatical.
Model Controlled Grammatical Perturbation on CLOTH
Original
accuracy
Perturbed
accuracy
Consistency
Backbone 76.21 72.34 91.90
LoRA-only 78.47 74.76 91.05
Direct-Joint 80.11 79.02 94.18
GTCA 81.68 80.42 98.27
Appendix Table 4: Performance under controlled grammatical perturbations on 1000 randomly sampled CLOTH questions. Perturbations preserve the gold answer while altering surface syntax. Consistency denotes the fraction of items whose predicted option remains unchanged across the original and perturbed questions. All results are on Qwen-2.5-7B, and we set temperature=0 and disabled sampling (do_sample=False).
Model Controlled Surface Noise on MMLU
Clean
accuracy
Noisy
accuracy
Consistency
Backbone 67.50 64.80 82.67
LoRA-only 69.10 66.00 82.60
Direct-Joint 70.20 68.60 84.43
GTCA 72.00 71.10 88.95
Appendix Table 5: Performance under controlled surface noise on 1000 randomly sampled MMLU questions. Noise is applied only to the question region and preserves the intended meaning. Consistency denotes the fraction of items whose predicted option remains unchanged across the clean and noisy questions. All results are on Qwen-2.5-7B, and we set temperature=0 and disabled sampling (do_sample=False).
Appendix F Robustness under Controlled Perturbations and Surface Noise
We construct a controlled grammatical perturbation suite on 1000 randomly sampled CLOTH questions. We apply meaning preserving transformations to the question region, including active passive alternation, relative clause insertion, and adjunct reordering, while keeping the gold answer unchanged. As shown in Appendix Table 4, GTCA achieves the highest perturbed accuracy and the highest consistency. This indicates that GTCA reduces answer flips under syntactic variation and better preserves decision stability when surface form changes but meaning remains constant.
We further test robustness to meaning preserving surface noise, including punctuation deletion, mild typos, and whitespace corruption, on 1000 randomly sampled MMLU questions. Appendix Table 5 shows that GTCA also yields the highest noisy accuracy and the highest consistency under this setting. These results suggest that the structural pathway remains beneficial even when parser inputs are mildly degraded, although the method still relies on parser quality as discussed in the limitations.